flux quantum (fluxoid) is approximately equal to 2à ƒ ƒ ƒ —10à ƒ ƒ ¢ ˆ ’7 gaussà ƒ ƒ ¢ ˆ ’cm2 A type II superconductor is placed in a small magnetic field, which is then slowly increased till the field starts penetrating the superconductor. The strength of the field at this point is 2à ƒ ƒ  €Ãƒ ƒ ƒ ƒ —105 gauss

The penetration depth of this superconductor is
a. 100 à ƒ ƒ ƒ …
b. 200 à ƒ ƒ ƒ …
c. 1000 à ƒ ƒ ƒ …
d. 314 à ƒ ƒ ƒ …
1
Expert's answer
2021-10-18T11:12:25-0400

The equation that ties the first critical field and penetration depth:


"H_{c1}=\frac{\phi_0}{\pi\lambda^2}→\lambda=\sqrt\frac{\phi_0}{\pi H_{c1}},\\\space\\
\lambda=100\text Å."